# VersaTune: Harnessing Vertical Domain Insights for Multi-Ability LLM Supervised Fine-Tuning

### v1

@@ 3.1 Preliminary: Learnable Potential and Forgetting Degree of Domain Knowledge
Mastery Ceiling. We first fine-tuned the small reference model for epochs on each individual domain separately, and identified the epoch with the lowest average loss during this process as the lower bound on the minimum loss attainable by the target model for the given domain. This value represents the highest level of domain knowledge mastery that the target model can achieve in the context of the current specific domain under given conditions.
Learnable Potential. We can observe whether a domain could be effectively learned by the model through comparing the difference between the loss of the target model and the minimum loss that the reference model can achieve. Based on these principles, we propose Equation 2 to score the learnable potential of domain .
@@
3.2 Effective Multi-Ability Fostering
@@
3.3 Flexible Domain Expansion
@@
When conducting fine-tuning on a pre-trained model, there are instances where we aim to particularly enhance the model’s performance on specific domain tasks. Consequently, our algorithmic framework ought to possess the flexibility to accommodate domain expansion and generalize effectively. Building upon Statement 1, we present the following statement tailored for domain expansion:
@@
As outlined in Algorithm 3, we initially establish the data distribution based on the knowledge detected from the original pre-trained model (Line 1). At each training step , we calculate the learnable potential and forgetting degree scores for each domain (Line 4-5), and assign domain weights for the current phase following the method from Algorithm 2 (Line 6). A trade-off is necessary between the remaining learning margin of the domain that requires focused cultivation and the model’s forgetting degree towards other domains: if the improvement benefit of the specific domain exceeds the average forgetting degree of the other domains (ratio greater than ), we increase the data weight of the current specific domain by , and proportionally reduce the weights of the other domains according to Equation 5 (Line 8-9); otherwise, we maintain the current domain distribution and only perform minor adjustments and renormalization as described in Algorithm 2 (Line 10-11). Subsequently, we update the parameters of the target model (Line 13).
4 Experiments
In this section, we describe details of our experimental setup (Section 4.1), the baseline methods we use for comparison (Section 4.2), and experimental results (Section 4.3).
| Model | Method | Law | Medical | Finance | Science | Code | General | Avg. |
| LLaMA-2-7B | Uniform Distribution | 23.22 | 25.51 | 18.11 | 26.10 | 4.56 | 45.78 | 23.88 |
| Inverse Distribution | 20.09↓ | 23.95↓ | 17.75↓ | 24.33↓ | 2.95↓ | 45.08↓ | 22.36↓ | |
| VersaTune | 29.75↑ | 37.90↑ | 32.66↑ | 53.41↑ | 15.65↑ | 48.22↑ | 36.27↑ | |
| LLaMA-3-8B | Uniform Distribution | 32.34 | 20.65 | 19.90 | 28.34 | 25.90 | 48.57 | 29.28 |
| Inverse Distribution | 27.66↓ | 17.14↓ | 18.51↓ | 26.36↓ | 22.01↓ | 47.93↓ | 26.60↓ | |
| VersaTune | 43.77↑ | 43.97↑ | 41.19↑ | 58.64↑ | 32.28↑ | 51.20↑ | 45.18↑ | |
| Qwen-1.5-7B | Uniform Distribution | 31.91 | 16.71 | 34.02 | 29.15 | 29.36 | 54.50 | 32.61 |
| Inverse Distribution | 29.97↓ | 13.03↓ | 33.12↓ | 27.53↓ | 26.25↓ | 54.58↓ | 30.75↓ | |
| VersaTune | 38.00↑ | 31.36↑ | 56.08↑ | 54.44↑ | 36.15↑ | 57.24↑ | 45.55↑ | |
| Qwen-2-7B | Uniform Distribution | 35.52 | 26.40 | 37.06 | 44.02 | 45.54 | 60.46 | 41.50 |
| Inverse Distribution | 30.91↓ | 20.11↓ | 33.92↓ | 38.41↓ | 42.14↓ | 59.63↓ | 37.52↓ | |
| VersaTune | 43.05↑ | 43.36↑ | 64.67↑ | 61.00↑ | 52.90↑ | 60.85↑ | 54.31↑ | |
| Qwen-2.5-7B | Uniform Distribution | 35.80 | 25.51 | 45.62 | 42.86 | 51.25 | 59.56 | 43.43 |
| Inverse Distribution | 31.67↓ | 21.76↓ | 43.16↓ | 39.73↓ | 48.80↓ | 59.30↓ | 40.74↓ | |
| VersaTune | 44.20↑ | 43.19↑ | 65.79↑ | 71.15↑ | 60.97↑ | 61.82↑ | 57.85↑ | |
| LLaMA-2-13B | Uniform Distribution | 41.26 | 34.76 | 35.20 | 48.25 | 37.55 | 55.20 | 42.04 |
| Inverse Distribution | 35.40↓ | 27.53↓ | 31.72↓ | 43.82↓ | 32.31↓ | 54.89↓ | 37.61↓ | |
| VersaTune | 48.01↑ | 46.73↑ | 47.30↑ | 63.36↑ | 46.15↑ | 57.14↑ | 51.45↑ | |
| Qwen-2.5-14B | Uniform Distribution | 45.11 | 43.28 | 64.56 | 56.60 | 70.45 | 65.4 | 57.57 |
| Inverse Distribution | 40.72↓ | 39.16↓ | 62.71↓ | 47.95↓ | 65.93↓ | 65.15↓ | 53.60↓ | |
| VersaTune | 53.59↑ | 47.70↑ | 72.26↑ | 72.88↑ | 82.06↑ | 65.75↑ | 65.71↑ |
4.1 Experimental Setup
We have collected SFT datasets spanning 6 vertical domains, including Lawyer-Instruct111https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-Instruct, the training portion of MedQA (Jin et al., 2020), Finance Alpaca222https://huggingface.co/datasets/gbharti/finance-alpaca, Sonnet3.5 Science Conversations333https://huggingface.co/datasets/jeffmeloy/sonnet3.5_science_conversations, Code Alpaca444https://github.com/sahil280114/codealpaca and Alpaca (Taori et al., 2023), denoted as , to represent law, medical, finance, science, code and general capabilities SFT datasets respectively. In order to prevent domain overlap, we curated the Alpaca dataset by excluding data pertaining to the other five specific domains, keeping only the general domain instances unrelated to them.
We use LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b) and Qwen (Bai et al., 2023; Yang et al., 2024) series as our pretrained language models . During the fine-tuning procedure, we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 4 epochs. To maintain consistency, the total volume of training data across domains was controlled to 60,000 per epoch. We conducted all fine-tuning and evaluation experiments on NVIDIA RTX H800.
4.2 Baselines
We compare VersaTune with the following baselines. For the scenario of effective multi-ability fostering: (1) The simplest baseline is uniform distribution, where each domain has an equal weight proportion. (2) Inverse distribution assigns the proportionate weights to each domain in an inverse manner to the detected knowledge distribution. Under the case of flexible domain expansion: (1) 100% specific domain strategy is a common practice to employ datasets consisting exclusively of data from a single domain during the fine-tuning stage. (2) Domain increase with uniform distribution of remainder elevates the proportion of a specific domain, while the remaining domains receive the balance of the distribution in an evenly distributed manner.
4.3 Results
We conduct evaluations to validate the efficiency of VersaTune across different models in scenarios that encompass both effective multi-ability fostering and flexible domain expansion, the main results of which are presented in Table 1 and Figure 2. We summarize the observations below.
VersaTune is efficient across different models in both scenarios. Table 1 shows that VersaTune consistently outperforms other methods across different models in terms of domain-specific capabilities. Compared to the uniform distribution of data across domains, VersaTune enhances downstream task performances by , which further underscores the effectiveness of our data composition strategy for enhancing the model’s overall multi-domain capabilities during the fine-tuning phase. For domain expansion scenarios, VersaTune has nearly maintained training efficiency while reducing the model’s loss of competencies in other domains by comparing to 100% specific domain fine-tuning, as depicted in Figure 2. Detailed results and analysis can be found in Appendix D.
Knowledge consistency training boosts performance. In Table 1, we present the experimental results of data composition strategies that allocate domain data inversely proportional to the detected knowledge distribution. As expected, the inverse distribution strategy yielded lower performance compared to the simplest approach of uniform distribution across all domains. This finding underscores the importance of aligning domain data ratios with the inherent knowledge distribution of the model during training.
5 Ablations and Analysis
Previously in Section 4, we demonstrated the effectiveness of VersaTune in enhancing multiple abilities and enabling flexible domain expansion during the model’s fine-tuning phase. In this section, we perform an in-depth analysis of VersaTune, where we ablate the components of dynamic adaptation in Algorithm 2, the criteria for determining the upper limit of domain expansion in Algorithm 3, and the impact of reference model scale employed for gauging the learnable potential and forgetting degree of the target model.
Dynamic adjustment enhances the robustness. During the process of cultivating multiple capabilities, we compared VersaTune with fixed domain weights derived from the knowledge distribution obtained from probing the target model prior to fine-tuning. Table 4 and Figure 3 demonstrated the high robustness of VersaTune, which dynamically adjusts domain weights during training by monitoring the learnable potential of each domain. In contrast, training with fixed domain weights exhibits certain fluctuations. A key reason for this phenomenon is the distribution of domain knowledge mastered by the model changes throughout training, and the learning efficiency varies among domains. Therefore, dynamically adjusting domain data weights based on the model’s feedback at different stages of training is crucial.
Establishing proportion thresholds for specific domains counts during domain expansion. We consider conducting a comparative analysis between the outcomes of VersaTune and those implementing an unconditional dynamic increase of the specific domain, where we remove the implementation of Line 8 in Algorithm 3. Figure 4 shows that the criteria for determining the upper limit on the proportion of a specific domain during domain expansion, has mitigated the loss of capabilities in other domains experienced by the target model during the fine-tuning process. Concurrently, it ensures gains in the capacity for the current domain of interest. We speculate that it is because by the later stages of fine-tuning, models’ proficiency in the target domain approaches saturation. Further increasing the proportion of the current domain provides diminishing returns and can lead to a significant loss of performance in other domains.
6 Related Work
Supervised Fine-Tuning of LLMs. The Supervised Fine-Tuning (SFT) stage of a Large Language Model (LLM) involves further training to refine the model’s task-solving capabilities and ensure greater alignment with human instructions (Zhao et al., 2023a). While recent research has delved into exploring fine-tuning methods for multi-task enhancement (Dong et al., 2023; Sanh et al., 2021), they are still in their early stages. However, as shown by proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), which exhibit outstanding multi-task performance, improving a model’s versatile capabilities across various domains during the SFT phase is crucial. Therefore, our work systematically investigates methods to enhance multi-domain performance during the SFT stage.
Data Reweighting for LLM Training. Data reweighting maintains full access to the entire dataset while adjusts the relative importance of each instance for various target tasks, which is essential for both pretraining and fine-tuning stages of LLMs (Wang et al., 2023). During the pretraining stage, DoReMi (Xie et al., 2024b) and DoGE (Fan et al., 2023) employ lightweight proxy models to estimate weights for different data sources, which are subsequently applied to the formal training of LLMs. Furthermore, Sheared LLaMA (Xia et al., 2023) implements an online variant of DoReMi, where the loss reference provided by the proxy model is replaced with the state of the pre-trained model from the previous training step. Additionally, ODM (Albalak et al., 2023) leverages the differences in loss or perplexity across domains within a multi-armed bandit framework to dynamically adjust the data distribution along training steps. As for the SFT phase, Dong et al. (2023) focus on enhancing the model’s math reasoning, coding, and general human-aligning abilities through a dual-stage mixed fine-tuning strategy. However, the mixing ratios for different domains rely heavily on enumeration and empirical methods, and the covered domains are not holistic. This paper provides a comprehensive overview of the model’s capabilities across various domains during the SFT stage and proposes appropriate multi-ability fine-tuning methods.
Knowledge Detection in LLMs. Investigating the knowledge contained in current LLMs is essential for guiding their subsequent training (Chang et al., 2024). The knowledge encompasses multiple dimensions, such as different domain sources and task attributes. Existing work in LLM knowledge detection primarily focuses on prompting and calibration. Directly prompting the model to generate sequences and extracting confidence scores from the model (Gekhman et al., 2024; Kadavath et al., 2022; Kuhn et al., 2023; Manakul et al., 2023) is a common strategy. However, such approaches are highly dependent on prompt design and task selection, which can introduce bias into the assessment. Other studies have attempted to infer the training data mixtures used in previous training stages of LLMs (Antoniades et al., 2024; Hayase et al., 2024; Hu et al., 2022; Ye et al., 2022). The essence of these studies is to evaluate the current knowledge state of the models and provide targeted strategies for data organization in subsequent training phases.
7 Conclusion
The versatile capability of Large Language Models (LLMs) across multiple domain tasks is of paramount importance. Existing fine-tuning methods for LLMs often focus on single domain tasks, leading to significant degradation in performance across other domains. In this paper, we introduce VersaTune, a data composition framework designed to enhance the multi-domain capabilities of models during the fine-tuning phase, based on the domain knowledge distribution of the target model. Experimental results from different models and downstream tasks across various domains demonstrate that VersaTune achieves excellent training outcomes in both scenarios of overall multi-domain enhancement and flexible expansion of specific domains.
Appendix A Details of Knowledge Distribution Detection
@@
During the domain knowledge detection phase for our target models, we employed Qwen2.5-72B-Instruct555https://huggingface.co/Qwen/Qwen2.5-72B-Instruct as the proprietary model . For each target model slated for fine-tuning, we prompted the generation of data samples using the Beginning of Sequence () token, with the sample number set at . These samples were subsequently assessed by the proprietary model to ascertain their probabilistic affinity for several vertical domains, including law, medicine, finance, science, code, and others. To ensure the reliability of our statistical outcomes, the entire process was iterated 5 times, with the maximum number of iterations set at . The average knowledge distribution was then computed across these iterations. Empirically, with a dataset of samples, the distribution of sequences generated by across domains demonstrated a high degree of consistency, with an overall variance not exceeding 1.874%. The final domain knowledge distribution for each open-source model is depicted in the stacked bar chart presented in Figure 5.
Appendix B Training
B.1 Training Datasets
For training, we selected representative datasets for each domain, which exhibit significant differences in format, sentence length, and domain-specific content. These differences reflect the heterogeneity of training data across various domains during the fine-tuning stage. Further details about these datasets can be found in Table 2. Specifically, for the Alpaca dataset, which we utilize for representing the general domain, we have excluded data related to law, medicine, finance, science, and code domains to ensure the precision and authenticity of the actual domain weight.
@@ B.1 Training Datasets
| Lawyer-Instruct | 9241 | Reformatted from LawyerChat Dataset666https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-chat | 1 | |
@@
B.2 Training Details
All experiments were conducted based on full fine-tuning. We employed a learning rate scheduler that incorporates linear warm-up and cosine decay, peaking at a learning rate of 2e-5. A batch size of 128 was utilized, and the training spanned 4 epochs across all selected datasets. For scenarios aimed at fostering multi-ability, we trained and assessed models including LLaMA-2-7B, LLaMA-3-8B, LLaMA-2-13B, Qwen-1.5-7B, Qwen-2-7B, Qwen-2.5-7B, and Qwen-2.5-14B. In the context of domain expansion, the training and evaluations were performed using the LLaMA-3-8B and LLaMA-2-13B models. The total number of samples per epoch was set to , with each domain’s samples being downsampled or upsampled according to the corresponding weights during the mixing process. Regarding reference models, for the LLaMA series, we used the Sheared-LLaMA-1.3B (Xia et al., 2023) as a lightweight reference model; as for the Qwen series, we utilized Qwen-2-1.5B and Qwen-2.5-1.5B as our reference models.
Appendix C Evaluation Details
We evaluate the performance of the models on the downstream tasks across various domains, using two relevant benchmarks for each domain. Details of the datasets are provided in Table 3. Specifically, for the MedMCQA dataset, since the standard answers for the test set are not publicly available, we conducted our evaluations using the validation dataset. For the MMLU dataset, we selected 14 sub-tasks to construct the MMLU-Sci subset (Zhang et al., 2024) for testing.
@@ Appendix C Evaluation Details
| FinanceIQ (Zhang & Yang, 2023) | 7,173 (10 sub-tasks) | Chinese | 5 | |

### v3

@@ 3.2.1. Learnable Potential and Forgetting Degree of Knowledge
Mastery Ceiling. We first fine-tuned the small reference model for epochs on each domain separately, and identified the epoch with the lowest average loss during this process as the lower bound on the minimum loss attainable by the target model for the given domain. This value represents the highest level of domain knowledge mastery that the model can achieve in the context of the current specific domain under given conditions.
Learnable Potential. We can observe whether a domain could be effectively learned by the model through comparing the difference between the loss of the target model and the minimum loss that the reference model can achieve. Based on these principles, we propose Equation 3 to score the learnable potential of domain :
@@
3.2.2. Effective Multi-Ability Fostering
@@
3.2.3. Flexible Domain Expansion
@@
When conducting fine-tuning on a pretrained model, there are instances where we aim to particularly enhance the model’s performance on specific domain tasks. Consequently, our algorithmic framework ought to possess the flexibility to accommodate domain expansion and generalize effectively. Building upon Statement 1, we present the following statement tailored for domain expansion:
@@
As outlined in Algorithm 3, we initially establish the data distribution based on the knowledge detected from the original pretrained model (Line 1). At each training step , we calculate the learnable potential and forgetting degree scores for each domain (Line 4-5), and assign domain weights for the current training phase following the method from Algorithm 2 (Line 6). A trade-off is necessary between the remaining learning margin of the domain that requires focused cultivation and the model’s forgetting degree towards other non-target domains: if the improvement benefit of the specific domain exceeds the average forgetting degree of the other domains (ratio greater than ), we increase the data weight of the current specific domain by , and proportionally reduce the weights of the other non-target domains according to Equation 6 (Line 8-9); otherwise, we maintain the current domain distribution and only perform minor adjustments and renormalization as described in Algorithm 2 (Line 10-11). Subsequently, we update the parameters of the target model (Line 13).
4. Implementation
VersaTune offers a comprehensive solution for data composition aimed at enhancing performance across multiple domains. In this section, we will briefly discuss the implementation of key components and the setting of hyperparameters.
@@ 4. Implementation
Domain Knowledge Detection. During the knowledge distribution detection phase for our target models, we employed Qwen2.5-72B-Instruct111https://huggingface.co/Qwen/Qwen2.5-72B-Instruct as the proprietary model . For each target model slated for fine-tuning, we prompted the generation of data samples using the Beginning of Sequence () token, with the sample number set at . These samples were subsequently assessed by the proprietary model to ascertain their probabilistic affinity for several domains, including law, medicine, finance, science, code, and others. To ensure the reliability of our statistical outcomes, the entire process was iterated 5 times, with the maximum number of iterations set at . The average knowledge distribution was then computed across these iterations. Empirically, with a dataset of samples, the distribution of sequences generated by across domains demonstrated a high degree of consistency, with an overall variance not exceeding 1.874%. The final domain knowledge distribution for each open-source model is depicted in the stacked bar chart presented in Figure 4. The pre-existing domain knowledge distribution varies among different models. Therefore, it is essential to develop a data composition strategy that is tailored to the specific model being trained.
Hyper-Parameters Setting. During the multi-domain task fine-tuning of LLMs, we configured the number of training steps in Algorithm 2 and Algorithm 3 to 4 epochs. We experimented with various magnitude of adjustment, specifically , and observed consistent weight ordering across domains, which far outperformed our baselines (detailed in Section 5.2). Based on these experimental outcomes, we set the magnitude of adjustment to . Additionally, in the context of domain expansion, we set the increment for the target domain to per training step, considering the overall domain weight distribution across models. The variation threshold, denoted as , reflects the trade-off between enhancing specific domain skills and mitigating the loss of capabilities in non-target domains, where we assigned a weight of .
5. Experiments
@@
| Lawyer-Instruct | 9241 | Reformatted from LawyerChat Dataset222https://huggingface.co/datasets/Alignment-Lab-AI/Lawyer-chat | 1 | |
@@
| FinanceIQ (Zhang and Yang, 2023) | 7,173 (10 sub-tasks) | Chinese | 5 | |